### v2

# OminiControl: Minimal and Universal Control for Diffusion Transformer

Abstract
In this paper, we introduce OminiControl, a highly versatile and parameter-efficient framework that integrates image conditions into pre-trained Diffusion Transformer (DiT) models. At its core, OminiControl leverages a parameter reuse mechanism, enabling the DiT to encode image conditions using itself as a powerful backbone and process them with its flexible multi-modal attention processors. Unlike existing methods, which rely heavily on additional encoder modules with complex architectures, OminiControl (1) effectively and efficiently incorporates injected image conditions with only 0.10.1~{}0.10.1additional parameters, and (2) addresses a wide range of image conditioning tasks in a unified manner, including subject-driven generation and spatially-aligned conditions such as edges, depth, and more. Remarkably, these capabilities are achieved by training on images generated by the DiT itself, which is particularly beneficial for subject-driven generation. Extensive evaluations demonstrate that OminiControl outperforms existing UNet-based and DiT-adapted models in both subject-driven and spatially-aligned conditional generation. Additionally, we release our training dataset, Subjects200K, a diverse collection of over 200,000 identity-consistent images, along with an efficient data synthesis pipeline to advance research in subject-consistent generation.  [^1] Code and dataset are available at https://github.com/Yuanshi9815/OminiControl
Figure 1: Results of our OminiControl on both subject-driven generation (top) and spatially-aligned tasks (bottom). The small images in red boxes show the input conditions.
1 Introduction
Diffusion models[9, 28, 25] have revolutionized the field of visual generation, demonstrating remarkable capabilities that significantly outperform traditional approaches like Generative Adversarial Networks (GANs)[6] in terms of image quality and diversity. While these models excel at generating highly realistic imagery, a critical challenge persists: enabling precise and flexible control over the generation process to accommodate diverse and complex user requirements.
Text-based conditioning has been a cornerstone in advancing controllable generation[28, 25, 13, 2, 23, 35], offering an intuitive interface for users to specify their desired outputs. However, text prompts alone often fail to convey precise spatial details and structural attributes that users wish to control. Consequently, recent research has explored complementary conditioning modalities for guiding diffusion models, with image-based control emerging as a particularly effective approach[41, 39, 22, 43, 15]. This multi-modal conditioning strategy enables more detailed and accurate control over the generation process, addressing the limitations inherent in purely text-based interfaces.
Current image conditioning methods can be broadly categorized into spatially aligned and non-spatially aligned approaches. Spatially aligned tasks such as sketch-to-image and inpainting require direct correspondence between conditioning and output images, typically achieved through methods like ControlNet[41] that inject conditioning features in a spatially-preserving manner. In contrast, non-spatially aligned applications including subject driven generation and style transfer, as demonstrated by IP-Adapter[39], often employ pre-trained encoders like CLIP[27] to extract global features for integration via cross-attention mechanisms.
Despite the effectiveness of existing image-conditioned approaches, they present several limitations that hinder their efficiency and flexibility [41, 39, 22]. First, most existing methods are designed specifically for UNet-based architectures [29, 41, 22, 39, 31, 42, 43, 40, 24, 16], as seen in Stable Diffusion models [28, 25]. While these approaches work well with UNet’s encoder-decoder structure, they may not translate effectively to the more advanced Diffusion Transformer (DiT) models[23] that have demonstrated superior image generation quality[13, 2]. Additionally, current approaches typically specialize in either spatially aligned [22, 43, 41] or non-spatially aligned tasks [39, 42, 17, 15, 12], lacking a unified architecture to handle both control types effectively. This specialization often requires practitioners to employ different methods for different control scenarios, increasing system complexity and implementation overhead. Furthermore, these methods rely heavily on additional network structures [39, 42, 22, 43, 41, 17], which introduce substantial parameter overhead.
To address these limitations, we propose a parameter-efficient approach for incorporating image-based control into DiT architectures[23]. Our method reuse the model’s existing VAE encoder[28] to process conditioning images. Following the same token processing pipeline as noisy image tokens, we augment the encoded features with learnable position embeddings[34], and seamlessly integrate them alongside latent noise in the denoising network. This design enables direct multi-modal attention interactions[23, 30] between condition and generation tokens throughout the DiT’s transformer blocks, facilitating efficient information exchange and control signal propagation.
Figure 2: Overview of the Diffusion Transformer (DiT) architecture and integration methods for image conditioning.
We implemented our method on the high-performing DiT-structured diffusion model, FLUX.1-dev[13], a large-scale model containing 12 billion parameters. Extensive experiments on edge-guided generation, depth-aware synthesis, region-specific editing, and identity-preserving generation indicate that our DiT-based approach yields better results compared to both UNet-based implementations[41, 39, 7] and their community adaptations on the FLUX.1 model [14, 37].
For identity-preserving generation, we developed a novel data synthesis pipeline that generates high-quality, identity-consistent image pairs. Using this pipeline, we created a comprehensive dataset comprising over 200,000 diverse images. To facilitate future research in this direction, we will release both our dataset and the complete pipeline implementation as open-source resources [^2] More details are available in the supplementary material..
In summary, we highlight our contributions as follows:
  1. 1.
    We present a parameter-efficient method for enabling image-conditioned control in Diffusion Transformer (DiT) models, achieving both spatially aligned and non-spatially aligned control within a unified framework.
  2. 2.
    We demonstrate the effectiveness of our approach through extensive experiments across diverse control tasks, including edge-guided generation, depth-aware synthesis, region-specific editing, and identity-preserving generation, consistently outperforming existing methods on both UNet implementations and their DiT adaptations.
  3. 3.
    We develop and release Subjects200K, a high-quality dataset of over 200,000 subject-consistent images, along with an efficient data synthesis pipeline, providing valuable resources to the research community for further exploration of subject-consistent generation tasks.
2 Related works
2.1 Diffusion-based models
Diffusion-based methods have emerged as a powerful framework for image generation[9, 28], demonstrating success across diverse tasks including text-to-image synthesis [28, 2, 33], image-to-image translation [32], and image editing [20, 1]. Recent advances have led to significant improvements in both quality and efficiency, notably through the introduction of latent diffusion models [28]. To further enhance generative capabilities, large-scale transformer architectures have been integrated into these frameworks, leading to advanced models like DiT[23, 2, 13, 3]. Building on these architectural innovations, FLUX[13] incorporates transformer-based design with flow matching objectives [18], achieving state-of-the-art generation performance.
2.2 Controllable generation with diffusion models
Controllable generation has been extensively studied in the context of diffusion models. Text-to-image models[28, 25] have established a foundation for conditional generation, while various approaches have been developed to incorporate additional control signals such as image. Notable methods include ControlNet [41], enabling spatially aligned control in diffusion models, and T2I-Adapter [22], which improves efficiency with lightweight adapters. UniControl [26] uses Mixture-of-Experts (MoE) to unify different spatial conditions, further reducing model size. However, these methods rely on spatially adding condition information to the denoising network’s hidden states, inherently limiting their effectiveness for non-spatial tasks like subject-driven generation. IP-Adapter [39] addresses this by introducing cross-attention through an additional encoder, and SSR-Encoder [42] further enhances identity preservation in image-conditioned tasks. Despite these advances [15, 19, 5], a unified solution for both spatially aligned and non-aligned tasks remains elusive.
3 Methods
3.1 Preliminary
The Diffusion Transformer (DiT) model [23], employed in architectures like FLUX.1 [13], Stable Diffusion 3 [28], and PixArt [2], uses a denoising network of transformer blocks to refine noisy image tokens iteratively.
Each transformer block processes two types of tokens: noisy image tokens $\mathbf{X}\in\mathbb{R}^{N\times d}$ and text condition tokens $\mathbf{C}_{\text{T}}\in\mathbb{R}^{M\times d}$, where $d$ is the embedding dimension, $N$ and $M$ are the number of image and text tokens respectively (Figure 2). These tokens are embedded into hidden states $X$ and $C_{\text{T}}$, which maintain consistent shapes throughout the transformer blocks.
In each DiT block, after normalizing $X$ and $C_{\text{T}}$, they are processed by the core MM-Attention module [30], which employs Rotary Position Embedding (RoPE) [34] to incorporate positional dependencies across tokens. For a token at position $(i,j)$ in the 2D grid, RoPE applies rotation matrices to the query and key projections:
$Q_{X}(i,j)$ $=W_{Q}\left(X_{i,j}\cdot R(i,j)\right),$ (1)
$K_{X}(i,j)$ $=W_{K}\left(X_{i,j}\cdot R(i,j)\right),$ (2)
where $R(i,j)$ is the rotation matrix at position $(i,j)$. Similarly, the text condition tokens $C_{\text{T}}$ have their query and key projections defined in the same way, with all text token positions set to $(0,0)$ in FLUX.1.
After applying RoPE, the queries, keys, and values from both token types are concatenated to form unified matrices $Q_{Z}$, $K_{Z}$, and $V_{Z}$, representing the combined token set $Z=[X;C_{\text{T}}]$. The MM-Attention operation is then computed as:
$\text{MMAttention}(Z)=\text{softmax}\left(\frac{Q_{Z}K_{Z}^{\top}}{\sqrt{d}}\right)V_{Z},$ (3)
enabling interactions between image and text condition tokens through the attention mechanism.
Figure 3: Comparison of results using two methods for integrating image conditions. The multi-modal approach demonstrates better condition following compared to direct addition.
3.2 Image condition integration
Our approach first encodes the condition image through the model’s VAE, projecting it into the same latent space as the noisy image tokens to form $\mathbf{C}_{\text{I}}\in\mathbb{R}^{N\times d}$.
Previous methods like ControlNet [41] and T2I-Adapter [22] incorporate the condition image by spatially aligning and adding its hidden states directly to those of the noisy image tokens:
$H_{X}=H_{X}+H_{C_{\text{I}}},$ (4)
where $H_{X}$ represents the combined hidden states for further processing, with $H_{C_{\text{I}}}$ being the hidden states from the condition image. While this approach proves effective for spatially aligned tasks, it faces two key limitations: (1) it lacks flexibility when handling non-aligned scenarios, and (2) even in spatially aligned cases, the direct addition of hidden states constrains token interactions, potentially limiting the model’s performance.
(a) Training losses for different image condition integration methods.
(b) Training loss for shared vs. shifting position.
Figure 4: Training loss comparisons.
In contrast, to enable non-aligned control tasks and provide greater conditioning flexibility, our method processes condition image tokens uniformly with text and noisy image tokens, integrating them into a unified sequence:
$Z=[X;C_{\text{T}};C_{\text{I}}],$ (5)
where $Z$ represents the concatenated sequence of noisy image tokens $X$, text tokens $C_{\text{T}}$, and condition image tokens $C_{\text{I}}$. This unified approach enables direct participation in multi-modal attention [30] without specialized processing pathways (illustrated in Figure 2).
Figure 5: (a) Attention maps for the Canny-to-image task, showing interactions between noisy image tokens $X$ and image condition tokens $C_{I}$. Strong diagonal patterns indicate effective spatial alignment. (b) Subject-driven generation task, with input condition and output image. Attention maps for $X\to C_{i}$ and $C_{i}\to X$ illustrate accurate subject-focused attention.
The comparative result shows that our approach achieves higher generation quality and better alignment with the conditions compared to the direct adding method, as illustrated in Figure 3. Moreover, the training curves in Figure 4(a) demonstrate that multi-modal attention method consistently achieves lower loss values than the direct adding approach. Besides, the effectiveness of this unified sequence approach is demonstrated across both spatially aligned and non-spatially aligned tasks (Figure 5), highlighting its versatility in handling diverse conditional generation scenarios.
3.3 Adaptive position embedding
Our unified sequence design allows for flexible integration of condition image tokens, but this requires incorporating positional information to ensure effective interaction with noisy image tokens. The relative positioning of these tokens is critical, as it directly affects the model’s learning efficiency and generalization capability.
In FLUX.1’s Transformers, each token is assigned a corresponding position index to encode spatial information. For a 512$\times$512 target image, the VAE [11] encoder first projects it into the latent space, then the latent representation is divided into a 32$\times$32 grid of tokens, where each token is assigned a unique two-dimensional position index $(i,j)$ with $i,j\in[0,31]$. This indexing scheme preserves the spatial structure of the original image in the latent space, while text tokens maintain a fixed position index of $(0,0)$.
For spatially aligned tasks, our initial approach was to assign condition tokens the same position embeddings as their corresponding tokens in the noisy image. However, for non-spatially aligned tasks such as subject-driven generation, our experiments revealed that shifting the position indices of condition tokens leads to faster convergence (Figure 4(b)). Specifically, we shift the condition image tokens to indices $(i,j)$ where $i\in[0,31]$ and $j\in[32,64]$, ensuring no spatial overlap with the original image tokens $X$.
3.4 Condition strength factor
The unified attention mechanism we adopt not only enables flexible token interaction but also allows us to precisely control the influence of condition images. Specifically, we designed a method that allows for manual adjustment of the condition image’s effect during inference. For a given strength factor $\gamma$, setting $\gamma=0$ removes the condition image’s influence, resulting in an output based purely on the original input. At $\gamma=1$, the output fully reflects the condition image, and as $\gamma$ increases beyond 1, the condition’s effect becomes even more pronounced.
To achieve this controllability, we introduce a bias term into the original MM-Attention operation. Specifically, we modify Equation 3 to:
$\text{BiasedAttention}(Z)=\text{softmax}\left(\frac{Q_{Z}K_{Z}^{\top}}{\sqrt{d}}+\text{bias}(\gamma)\right)V_{Z},$ (6)
[…]
$\text{bias}(\gamma)=\begin{bmatrix}\mathbf{0}_{M\times M}&\mathbf{0}_{M\times N}&\mathbf{0}_{M\times N}\\ \mathbf{0}_{N\times M}&\mathbf{0}_{N\times N}&\log(\gamma)\mathbf{1}_{N\times N}\\ \mathbf{0}_{N\times M}&\log(\gamma)\mathbf{1}_{N\times N}&\mathbf{0}_{N\times N}\end{bmatrix}.$ (7)
This design ensures that the strength factor $\gamma$ only affects the attention weights between noisy image tokens and condition image tokens, while maintaining the original attention patterns for text tokens and within-modality interactions.
3.5 Subjects200K datasets
Figure 6: Examples from our Subjects200Kdataset. Each pair of images shows the same object in varying positions, angles, and lighting conditions. The dataset includes diverse objects such as clothing, furniture, vehicles, and animals, totaling over 200,000 images. This dataset, along with the generation pipeline, will be publicly released.
Training models for subject-consistent generation typically requires paired images that maintain identity consistency while exhibiting variations in pose, lighting, and other attributes. Previous methods like IP-Adapter [39] use identical images for conditioning and target pairs, which proves effective for their approaches. However, in our framework, this setup leads to overfitting, causing the model to generate outputs nearly identical to the inputs.
To overcome these limitations, we developed a dataset featuring images that preserve subject identity while incorporating natural variations. While existing datasets [31, 12, 15, 17] address similar needs, they often face constraints in either quality or scale. We therefore propose a novel synthesis pipeline leveraging FLUX’s inherent capability to generate pairs of visually related images from carefully crafted prompts.
Our pipeline utilizes ChatGPT-4o to generate over 20,000 diverse image descriptions, which guide FLUX in producing more than 200,000 images (Figure 6). The generated images undergo quality assessment using ChatGPT-4o’s vision capabilities, ensuring both consistency and diversity in the final dataset. The curated dataset and complete generation pipeline are publicly available [^3] More details are provided in the supplementary material. Dataset and code are available at https://github.com/Yuanshi9815/Subjects200K.
4 Experiment
4.1 Setup
Tasks and base model. We evaluate our method on two categories of conditional generation tasks: spatially aligned tasks (including Canny-to-image, depth-to-image, masked-based inpainting, and colorization) and subject-driven generation. We build our method upon FLUX.1 [13], a latent rectified flow transformer model for image generation. By default, we use FLUX.1-dev to generate images for spatially aligned tasks. In subject-driven generation tasks, we switch to FLUX.1-schnell as we observed it tend to produce better visual quality.
Condition Model Method Controllability General Quality Text Consistency
[…]
Canny SD1.5 ControlNet 0.34 18.74 0.35 0.45 67.81 0.75
T2I-Adapter 0.22 20.06 0.35 0.39 67.88 0.74
FLUX.1 ControlNet 0.21 98.68 0.25 0.37 56.90 0.53
Ours 0.38 20.63 0.40 0.61 75.91 0.76
Depth SD1.5 ControlNet 923 23.02 0.34 0.47 70.73 0.726
T2I-Adapter 1560 24.72 0.27 0.39 69.99 0.72
FLUX.1 ControlNet 2958 62.20 0.26 0.38 66.84 0.54
Ours 903 27.26 0.39 0.55 75.06 0.728
Deblur FLUX.1 ControlNet 572 30.38 0.74 0.31 54.37 0.78
Ours 132 11.49 0.87 0.39 67.63 0.87
Colorization FLUX.1 ControlNet 351 16.27 0.64 0.43 70.95 0.85
Ours 24 10.23 0.73 0.43 70.74 0.90
Mask SD1.5 ControlNet 7588 13.14 0.40 0.41 67.22 0.84
FLUX.1 Ours 6248 15.66 0.48 0.45 72.61 0.80
Table 1: Quantitative comparison with baseline methods on five spatially aligned tasks. We evaluate methods based on Controllability (F1-Score for Canny, MSE for others), General Quality (FID, SSIM, MAN-IQA, MUSIQ), and Text Consistency (CLIP-Score). For F1-Score, higher is better; for MSE, lower is better. Best results are shown in bold.
Implement details. Our method utilizes LoRA [4]for fine-tuning the base model with a default rank of 4. To preserve the model’s original capabilities and achieve flexibility, the LoRA scale is set to 0 when processing non-condition tokens by default.
Training. Our model is trained with a batch size of 1 and gradient accumulation over 8 steps (effective batch size of 8). We employ the Prodigy optimizer [21] with safeguard warmup and bias correction enabled, setting the weight decay to 0.01. The experiments are conducted on 2 NVIDIA H100 GPUs (80GB each). For spatially aligned tasks, models are trained for 50,000 iterations, while subject-driven generation models are trained for 15,000 iterations.
Baselines. For spatially-aligned tasks, we compare our method with both the original ControlNet [41] and T2I-Adapter [22] on Stable Diffusion 1.5, as well as ControlNetPro [14], the FLUX.1 implementation of ControlNet. For subject-driven generation, we compare with IP-Adapter [39], evaluating its implementations FLUX.1 [37].
Evaluation metrics. We evaluate our model on both spatially aligned tasks and subject-driven generation. For spatially aligned tasks, we assess two aspects: generation quality and controllability. Generation quality is measured using FID [8], SSIM, MAN-IQA [38], and MUSIQ [10] for visual fidelity, along with CLIP Score [27] for semantic consistency. For controllability, we compute F1 Score between extracted and input edge maps in edge-conditioned generation, and MSE between extracted and original condition maps for other tasks (using Depth Anything for depth, color channel separation for colorization, etc.). For subject-driven generation, we propose a five-criteria framework evaluating both preservation of subject characteristics (identity preservation, material quality, color fidelity, natural appearance) and accuracy of requested modifications, with all assessments conducted through GPT-4o’s vision capabilities to ensure systematic evaluation. Detailed evaluation methodology is presented in Appendix B.1.
Evaluation protocol. We conducted evaluations on two datasets. For spatially-aligned tasks, we use COCO 2017 validation set (5,000 images) resized to 512$\times$512, using task-specific conditions and associated captions as prompts with a fixed seed of 42. For subject-driven generation, we test on 750 text-condition pairs (30 subjects × 25 prompts) from DreamBooth[31] dataset with 5 different seeds, using one selected image per subject as the condition.
4.2 Main result
Figure 7: Qualitative results comparing different methods. Left: Spatially aligned tasks across Canny, depth, out-painting, deblurring, colorization. Right: Subject-driven generation with beverage can, shoes and robot toy. Our method demonstrates superior controllability and visual quality across all tasks.
Spatially aligned tasks As shown in Table 1, we comprehensively evaluate our method against existing approaches on five spatially aligned tasks. Our method achieves the highest F1-Score of 0.38 on depth-to-image generation, significantly outperforming both SD1.5-based methods ControlNet [41] and T2I-Adapter [22], as well as FLUX.1-based ControlNetPro [14]. In terms of general quality metrics, our approach demonstrates consistent superiority across most tasks, showing notably better performance in SSIM [36], MAN-IQA [38], and MUSIQ [10] scores. For challenging tasks like deblurring and colorization, our method achieves substantial improvements: the MSE is reduced by 77% and 93% respectively compared to ControlNetPro, while the FID scores [8] improve from 30.38 to 11.49 for deblurring. The CLIP-Score metrics [27] indicate that our method maintains high text-image consistency across all tasks, suggesting effective preservation of semantic alignment while achieving better control and visual quality. As shown in Figure 7, our method produces sharper details and more faithful color reproduction in colorization tasks, while maintaining better structural fidelity in edge-guided generation and deblurring scenarios.
Figure 8: Radar charts visualization comparing our method (blue) with baselines across five evaluation metrics.
Subject driven generation Figure 8 presents a comprehensive comparison against existing baselines. Our method demonstrates superior performance, particularly in identity preservation and modification accuracy. Averaging over random seeds, we achieve 75.8% modification accuracy compared to IP-Adapter (FLUX)’s 57.7%, while maintaining 50.6% identity preservation against IP-Adapter (SD 1.5)’s 29.4%. The advantage amplifies in best-seed scenarios, achieving 90.7% modification accuracy and 82.3% identity preservation - surpassing the strongest baselines by 15.8 and 18.0 percentage points, demonstrating effective subject-fidelity editing. These quantitative results are further corroborated by user studies presented in Appendix B.1.
Comparative paramter efficiency. As shown in Table 2, our approach achieves remarkable parameter efficiency compared to existing methods. For the 12B parameter FLUX.1 model, our method requires only 14.5M trainable parameters (approximately 0.1%), which is significantly lower than ControlNet (27.5%) and IP-Adapter (7.6%). Even when utilizing the original VAE encoder from FLUX.1, our method still maintains high efficiency with just 0.4% additional parameters, demonstrating the effectiveness of our parameter-efficient design.
Methods Base model Parameters Ratio
ControlNet SD1.5 / 860M 361M $\sim$42%
T2I-Adapter 77M $\sim$9.0%
IP-Adapter 449M $\sim$52.2%
ControlNet FLUX.1 / 12B 3.3B $\sim$27.5%
IP-Adapter 918M $\sim$7.6%
Ours FLUX.1 / 12B 14.5M / 48.7M w/ Encoder $\sim$0.1/ $\sim$0.4w/ Encoder
Table 2: Additional parameters introduced by different image conditioning methods. For IP-Adapter, the parameter count includes the CLIP Image encoder. For our method, we also report results when using the original VAE encoder from FLUX.1.
4.3 Empirical studies
Effect of training data.
Figure 9: Comparison of models trained with different data. The model trained by data augmentation tends to copy inputs directly, while model trained by our Subjects200Kgenerates novel views while preserving identity.
For subject-driven generation, our model takes a reference image of a subject (e.g., a plush toy or an object) and a text description as input, aiming to generate novel images of the same subject following the text guidance while preserving its key characteristics.
To validate the effectiveness of our Subjects200K dataset described in Section 3.5, we compare two training strategies for this task. The first approach relies on traditional data augmentation, where we apply random cropping, rotation, scaling, and adjustments to contrast, saturation, and color to the original images. The second approach utilizes our Subjects200K dataset. As shown in Figure 9, the model trained with data augmentation only learns to replicate the input conditions with minimal changes. In the first row, it simply places the taco plush toy in a bright room setting while maintaining its exact appearance and pose. Similarly, in the second row, the yellow alarm clock is reproduced with nearly identical details despite the window-side placement instruction. In contrast, our Subjects200K-trained model demonstrates the ability to generate diverse yet consistent views of the subjects while faithfully following the text prompts.
Condition strength analysis. We evaluate our condition strength control through qualitative experiments. Figure 10 shows generated results with varying strength factor $\gamma\in{0.25,0.5,0.75,1.0}$. Results show that $\gamma$ effectively controls the generation process for both spatially-aligned tasks like depth-to-image generation and non-spatially-aligned tasks like subject-driven generation, enabling flexible control over the condition’s influence.
Figure 10: Comparison of models trained with different data. The model trained by data augmentation tends to copy inputs directly, while model trained by our Subjects200Kgenerates novel views while preserving identity.
Impact of LoRA rank. We conducted extensive experiments with different LoRA ranks (1, 2, 4, 8, and 16) for the Canny-to-image task. As shown in Table 3, our experiments show that increasing the LoRA rank generally improves model performance, with rank 16 achieving the best results across multiple aspects: image quality (measured by FID and SSIM), condition control capability (measured by F1 Score), while maintaining competitive text-image consistency (measured by CLIP-Score). Notably, even with smaller ranks (e.g., rank 1), the model demonstrates competitive performance, especially in text-image alignment where it achieves the highest CLIP-Score of 0.765, showing the efficiency of our approach even with limited parameters.
Study Setting FID $\downarrow$ SSIM $\uparrow$ F1 Score $\uparrow$ CLIP Score $\uparrow$
LoRA Rank 1 21.09 0.412 0.385 0.765
2 21.28 0.411 0.377 0.751
4 20.63 0.407 0.380 0.761
8 21.40 0.404 0.3881 0.761
16 19.71 0.425 0.407 0.764
Condition Blocks Early 25.66 0.369 0.23 0.72
Full 20.63 0.407 0.38 0.76
Table 3: Ablation studies on (1) LoRA rank for the Canny-to-image task and (2) condition signal integration approaches. Results show that LoRA rank of 16 and full-depth integration achieve the best performance. Rows with blue background indicate our default settings (LoRA rank=4, Full condition integration). Best results are in bold.
Conditioning depth. FLUX.1’s transformer architecture features two distinct types of blocks: early blocks that employ separate normalization modules for different modalities tokens (text and image) and later blocks that share unified normalization across all tokens. As shown in Table 3, experiments reveal that restricting condition signal integration to only these early blocks results in insufficient controllability over the generation process. This suggests that allowing the condition signals to influence the entire transformer stack is crucial for achieving the desired levels of control over the output. Notably, this finding indicates that the preview approaches[14, 37, 39, 41, 22] of inserting condition signals primarily in early blocks, which were effective in UNet-based architectures, may not fully translate to DiT-based models like FLUX.1.
5 Conclusion
OminiControl offers parameter-efficient image-conditioned control for Diffusion Transformers across diverse tasks using a unified token approach, without extra modules. Our method outperforms traditional approaches, and the new Subjects200K dataset—featuring over 200,000 high-quality, subject-consistent images—supports advancements in subject-consistent generation. Results confirm OminiControl’s scalability and effectiveness in diffusion models.
6 Acknowledgment
We would like to acknowledge that the computational work involved in this research work is partially supported by NUS IT’s Research Computing group using grant numbers NUSREC-HPC-00001.
Supplementary Material
Appendix A Details of Subjects200K datasets
We present a comprehensive synthetic dataset constructed to address the limitations in scale and image quality found in previous datasets [31, 12, 15, 17]. Our approach leverages FLUX.1-dev [13] to generate high-quality, consistent images of the same subject under various conditions.
Subjects200K dataset currently consists of two splits, both generated using similar pipelines. Split-1 contains paired images of objects in different scenes, while Split-2 pairs each object’s scene image with its corresponding studio photograph. Due to their methodological similarities, we primarily focus on describing the synthesis process and details of Split-2, although both splits are publicly available. Our complete Subjects200K dataset can be fully accessed via this link.
A.1 Generation pipeline
Our dataset generation process consists of three main stages: description generation, image synthesis, and quality assessment.
Description Generation We employed ChatGPT-4o to create a hierarchical structure of descriptions: We first generated 42 diverse object categories, including furniture, vehicles, electronics, clothing, and others. For each category, we created multiple object instances, totaling 4,696 unique objects. Each object entry consists of: (1) A brief description, (2) Eight diverse scene descriptions, (3) One studio photo description. Figure S2 shows a representative example of our structured description format.
Image Synthesis We designed a prompt template to leverage FLUX’s capability of generating paired images containing the same subject. Our template synthesizes a comprehensive prompt by combining a brief object description with two distinct scene descriptions, ensuring subject consistency while introducing environmental variations.
The detailed prompt structure is illustrated in Figure S3. For each prompt, we set the image dimensions to 1056×528 pixels and generated five images using different random seeds to ensure diversity in our dataset. During the training process, we first split the paired images horizontally, then performed central cropping to obtain 512×512 pixel image pairs. This padding strategy was implemented to address cases where the generated images were not precisely bisected, preventing potential artifacts from appearing in the wrong half of the split images.
Figure S1: Examples of successful and failed generation results from Subjects200K dataset. Green checks indicate successful cases where subject identity and characteristics are well preserved, while red crosses show failure cases.
Quality assessment We leveraged ChatGPT-4o’s vision capabilities to rigorously evaluate the quality of images generated by FLUX.1-dev. The assessment focused on multiple critical aspects:
  • Image composition: Verifying that each image properly contains two side-by-side views.
  • Subject consistency: Ensuring the subject maintains identity across both views.
  • Image quality: Confirming high resolution and visual fidelity.
To maintain stringent quality standards, each image underwent five independent evaluations by ChatGPT-4o. Only images that passed all five evaluations were included in our training dataset. Figure S1 presents representative examples from our quality-controlled dataset.
"brief_description":
"A finely-crafted wooden seating piece.",
"scene_descriptions": [
"Set on a sandy shore at dusk, it faces the ocean with a gentle breeze rustling nearby palms, bathed in soft, warm twilight.",
"Positioned in a bustling urban cafe, it stands out against exposed brick walls, capturing the midday sun through a wide bay window."
// Additional six scene descriptions omitted
"studio_photo_description":
"In a professional studio against a plain white backdrop, it is captured in three-quarter view under uniform high-key lighting, showcasing the delicate grain and smooth of its finely-crafted surfaces."
Figure S2: An example of our structured description format for dataset generation.
prompt_1 = f"Two side-by-side images of the same object: {brief_description}"
prompt_2 = f"Left: {scene_description1}"
prompt_3 = f"Right: {scene_description2}"
prompt_image = f"{prompt_1}; {prompt_2}; {prompt_3}"
Figure S3: Our prompt template for paired image generation. The template combines a brief object description with two distinct scene descriptions to maintain subject consistency while varying environmental conditions.
A.2 Dataset Statistics
In Split-2, we first generated 42 distinct object categories, from which we created and curated a set of 4,696 detailed object instances. Then we combine these descriptions to generate 211,320 subject-consistent image pairs. Through rigorous quality control using GPT-4o, we selected 111,767 high-quality image pairs for our final dataset. This extensive filtering process ensured the highest standards of image quality and subject consistency, resulting in a collection of 223,534 high-quality training images.
Appendix B Additional experimental results
B.1 Evaluation for subject-driven generation
Framework and criteria. To systematically evaluate subject-driven generation quality, we establish a framework with five criteria assessing both preservation of subject characteristics and accuracy of requested modifications:
  • Identity Preservation: Evaluates preservation of essential identifying features (e.g., logos, brand marks, distinctive patterns)
  • Material Quality: Assesses if material properties and surface characteristics are accurately represented
  • Color Fidelity: Evaluates if colors remain consistent in regions not specified for modification
  • Natural Appearance: Assesses if the generated image appears realistic and coherent
  • Modification Accuracy: Verifies if the changes specified in the text prompt are properly executed
Method Identity Material Color Natural Modification Average
preservation quality fidelity appearance accuracy score
Average over 5 random seeds
IP-Adapter (SD 1.5) 29.4 86.1 45.3 97.9 17.0 55.1
SSR-Encoder 46.0 92.0 54.2 96.3 28.5 63.4
IP-Adapter (FLUX) 11.8 65.8 30.8 98.1 57.7 52.8
Ours 50.6 84.3 55.0 98.5 75.8 72.8
Best score over 5 random seeds
IP-Adapter (SD 1.5) 56.3 98.9 70.1 99.7 37.2 72.5
SSR-Encoder 64.3 99.2 74.4 99.1 53.6 78.1
IP-Adapter (FLUX) 27.5 86.1 53.6 99.9 74.9 68.4
Ours 82.3 98.0 88.4 100.0 90.7 91.9
Table S1: Quantitative evaluation results (in percentage) across different evaluation criteria. Higher values indicate better performance.
Figure S4: User study results comparing different methods across three metrics: identity consistency, text-image alignment, and visual coherence.
User studies. To further validate our approach, we conducted user studies collecting 375 valid responses. Participants evaluated the generated images across three key dimensions: identity consistency, text-image alignment, and visual coherence between subjects and backgrounds. The results shown in Figure S4 corroborate our quantitative findings, with our method achieving superior performance across all evaluation criteria.
B.2 Additional generation results
We showcase more generation results from our method. Figure S5 presents additional results on the DreamBooth dataset, while Figure S6 demonstrates our method’s effectiveness on other subject-driven generation tasks.
Figure S5: More results on Dreambooth dataset.
Figure S6: More results on other subject-driven generation tasks.