### v4

# OminiControl: Minimal and Universal Control for Diffusion Transformer

Abstract
In this paper, we introduce OminiControl, a highly versatile and parameter-efficient framework that integrates image conditions into pre-trained Diffusion Transformer (DiT) models. At its core, OminiControl leverages a parameter reuse mechanism, enabling the DiT to encode image conditions using itself as a powerful backbone and process them with its flexible multi-modal attention processors. Unlike existing methods, which rely heavily on additional encoder modules with complex architectures, OminiControl (1) effectively and efficiently incorporates injected image conditions with only 0.10.1~{}0.10.1additional parameters, and (2) addresses a wide range of image conditioning tasks in a unified manner, including subject-driven generation and spatially-aligned conditions such as edges, depth, and more. Remarkably, these capabilities are achieved by training on images generated by the DiT itself, which is particularly beneficial for subject-driven generation. Extensive evaluations demonstrate that OminiControl outperforms existing UNet-based and DiT-adapted models in both subject-driven and spatially-aligned conditional generation. Additionally, we release our training dataset, Subjects200K, a diverse collection of over 200,000 identity-consistent images, along with an efficient data synthesis pipeline to advance research in subject-consistent generation.  [^1] Code and dataset are available at https://github.com/Yuanshi9815/OminiControl
Figure 1: Results of our OminiControl on both subject-driven generation (top) and spatially-aligned tasks (bottom). The small images in red boxes show the input conditions.
1 Introduction
The ability to generate high-quality images with precise user control remains a central challenge in computer vision. While diffusion models [10, 28] have significantly advanced the state of the art in image generation, surpassing traditional GAN-based approaches [7] in both quality and diversity, the problem of fine-grained control persists. Text-conditioned diffusion models [28, 25, 5, 14] serve as the primary paradigm for controllable generation, enabling users to guide the synthesis process through natural language descriptions. However, these models fundamentally lack the capacity to specify exact spatial details and visual attributes that users often need to control.
To address this limitation, recent works have turned to image-based control methods [42, 41, 22, 44, 16, 34, 38, 17, 26, 37] as a promising solution to achieve fine-grained control. These approaches enable users to specify their intentions through reference images or visual hints, offering more precise guidance than text alone [28].
However, current image control methods face several key challenges: First, existing methods require substantial architectural modifications with dedicated control modules, leading to significant parameter overhead [42, 41]. Second, these approaches show clear task bias – they typically work well for either spatially aligned controls (e.g., edge-guided or depth-guided generation [42, 22, 44, 26]) or spatially unaligned ones (e.g., style transfer or subject-driven generation [41, 17, 16, 34]), but rarely both. Third, current approaches are built primarily for UNet architectures [29]. These approaches may yield suboptimal results when applied to newer Diffusion Transformer (DiT) models [24] due to fundamental architectural differences, despite the latter’s superior generation capabilities [14, 2, 5].
These challenges, particularly within the emerging paradigm of Diffusion Transformer models, motivate us to rethink the fundamental approach to image control. We propose OminiControl, an omni-capable yet minimal framework that achieves effective control by leveraging existing components within the DiT architecture.
To achieve the goal of minimal architectural and parameters overhead, OminiControl first utilizes the built-in VAE encoder to project conditioning images into the same latent space as the target images. The conditional tokens are concatenated with the noisy latent and processed by the DiT using multi-modal attention [23, 30, 24]. This unified sequence allows the denoising network to handle both signals directly, requiring only minimal LoRA [4] fine-tuning of the existing transformer blocks. Unlike existing methods such as IP-Adapter [41] that relies on separate CLIP [27] feature extractors and ControlNet [15] that requires additional control modules, OminiControl significantly reduces architectural complexity and parameter overhead.
Building on this, OminiControl unifies controlled generation tasks using the DiT architecture. By modifying explicit position embeddings [35] in the attention mechanism, it handles both spatially aligned and non-aligned tasks. By explicitly controlling condition token positions in the attention mechanism, OminiControl not only handles non-aligned control tasks through semantic attention, but also achieves spatially-aligned control through flexible token interactions, rather than relying on rigid spatial feature addition as in ControlNet [42].
Alongside the design of OminiControl, to fully realize its potential in diverse control scenarios, we address a critical challenge in subject-driven generation – the scarcity of high-quality training data for identity preservation. To this end, we develop an automated data synthesis pipeline specifically designed for subject-driven generation tasks. This pipeline builds upon a key observation that DiT models can generate identity-consistent image pairs with remarkable consistency by simply providing appropriate prompts [14]. Using this pipeline, we create a comprehensive dataset Subjects200K including over 200,000 diverse images with rich identity variations. To facilitate future research, we have released both our dataset and the complete pipeline implementation as open-source resources [^2] Code and dataset are available at: https://github.com/Yuanshi9815/Subjects200K.
Extensive experiments demonstrate that our approach achieves better performance across diverse control tasks. Building upon the state-of-the-art DiT model Flux [14], we evaluate OminiControl on a comprehensive suite of tasks including edge-guided generation, depth-guided generation, and subject-driven generation. Our method outperforms existing approaches such as ControlNet [42], T2I-Adapter [22], and IP-Adapter [41] in their original UNet implementations, as well as their community adaptations on the DiT architecture [15, 39]. Notably, OminiControl achieves these superior results while maintaining remarkable efficiency – requiring only $0.1\%$ additional parameters compared to the base model.
In summary, we highlight our contributions as follows:
  • We propose OminiControl, a minimal and universal control framework for DiT models that: (1) minimizes architectural changes by reusing existing components , introducing only 0.1% additional parameters; and (2) achieves unified control by integrating condition tokens into the original multi-modal attention and leveraging position embeddings to handle both spatially-aligned and unaligned tasks.
  • We develop an automated data synthesis pipeline that leverages DiT’s inherent capability in generating identity-consistent image pairs. Using this, we create and release Subjects200K, a large-scale dataset containing over 200,000 diverse images for subject-driven generation.
  • We conduct extensive experiments on various control tasks including edge-guided, depth-guided, and subject-driven generation. Building upon Flux, OminiControl outperforms existing approaches in both their original UNet implementations and DiT adaptations.
Figure 2: Overview of the Diffusion Transformer (DiT) architecture and integration methods for image conditioning.
2 Related works
2.1 Diffusion models
Diffusion models have emerged as a powerful framework for image generation[10, 28], demonstrating success across diverse tasks including text-to-image synthesis [28, 2, 33], image-to-image translation [32], and image editing [20, 1]. Recent advances have led to significant improvements in both quality and efficiency, notably through the introduction of latent diffusion models [28]. To further enhance generative capabilities, large-scale transformer architectures have been integrated into these frameworks, leading to advanced models like DiT[24, 2, 14, 3]. Building on these architectural innovations, FLUX[14] incorporates transformer-based design with flow matching objectives [18], achieving state-of-the-art generation performance.
2.2 Controllable generation with diffusion models
Controllable generation has been extensively studied in the context of diffusion models. Text-to-image models[28, 25] have established a foundation for conditional generation, while various approaches have been developed to incorporate additional control signals such as image. Notable methods include ControlNet [42], enabling spatially aligned control in diffusion models, and T2I-Adapter [22], which improves efficiency with lightweight adapters. UniControl [26] uses Mixture-of-Experts (MoE) to unify different spatial conditions, further reducing model size. However, these methods rely on spatially adding condition features to the denoising network’s hidden states, inherently limiting their effectiveness for spatially non-aligned tasks like subject-driven generation. IP-Adapter [41] addresses this by introducing cross-attention through an additional encoder, and SSR-Encoder [43] further enhances identity preservation in image-conditioned tasks. Despite these advances [16, 19, 6], a unified solution for both spatially aligned and non-aligned tasks remains elusive.
3 Methods
3.1 Preliminary
The Diffusion Transformer (DiT) model [24], employed in architectures like FLUX.1 [14], Stable Diffusion 3 [28], and PixArt [2], uses transformer as denoising network to refine noisy image tokens iteratively.
A DiT model processes two types of tokens: noisy image tokens ${X}\in\mathbb{R}^{N\times d}$ and text condition tokens ${C}_{\text{T}}\in\mathbb{R}^{M\times d}$, where $d$ is the embedding dimension, $N$ and $M$ are the number of image and text tokens respectively (Figure 2). Throughout the network, these tokens maintain consistent shapes as they pass through multiple transformer blocks.
In FLUX.1, each DiT block consists of layer normalization followed by Multi-Modal Attention (MMA) [23], which incorporates Rotary Position Embedding (RoPE) [35] to encode spatial information. For image tokens ${X}$, RoPE applies rotation matrices based on the token’s position $(i,j)$ in the 2D grid:
${X}_{i,j}\rightarrow{X}_{i,j}\cdot R(i,j),$ (1)
where $R(i,j)$ is the rotation matrix at position $(i,j)$. Text tokens ${C}_{\text{T}}$ undergo the same transformation with their positions set to $(0,0)$.
The multi-modal attention mechanism then projects the position-encoded tokens into query $Q$, key $K$, and value $V$ representations. It enables the computation of attention between all tokens:
$\text{MMA}([{X};{C}_{\text{T}}])=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d}}\right)V,$ (2)
where $[{X};{C}_{\text{T}}]$ denotes the concatenation of image and text tokens. This formulation enables bidirectional attention.
Figure 3: Comparison of results using two methods for integrating image conditions. The multi-modal approach demonstrates better condition following compared to direct addition.
3.2 OminiControl
(a) Training losses for different image condition integration methods.
(b) Training loss for shared vs. shifting position.
Figure 4: Training loss comparisons.
Building upon the DiT architecture, with FLUX.1 [14] as our implementation base,we aim to develop OminiControl, an omni-capable yet minimal control framework that guides generation through visual conditions.
Leveraging existing architecture
To achieve both omni-capability and minimal architectural changes, OminiControl first reuses the VAE [12, 28] encoder from the base DiT model to project conditioning images into the same latent space as noisy image tokens. This approach contrasts with previous methods that rely on separate feature extractors like CLIP [27, 41] or additional control modules [43], significantly reducing architectural complexity.
The encoded condition tokens ${C}_{\text{I}}$ share the same dimensionality and latent space as the noisy image tokens ${X}$, enabling them to be directly processed by the transformer blocks. Since both condition and image tokens reside in the same latent space, OminiControl leverages the existing DiT blocks to process them jointly, requiring only minimal fine-tuning through LoRA [4] to adapt these blocks for handling the additional condition tokens, instead of duplicating the entire network as done in ControlNet [42].
Unified sequence processing
Building upon these shared latent space, we next need to integrate condition tokens into the model for flexible control.
Previous methods like ControlNet [42] and T2I-Adapter [22] incorporate condition images through direct feature addition:
${X}\leftarrow{X}+{C}_{\text{I}},$ (3)
where the condition features ${C}_{\text{I}}$ are spatially aligned and added to the noisy image tokens ${X}$. While effective for spatially aligned tasks, this approach faces two limitations: (1) it lacks flexibility for non-aligned scenarios where spatial correspondence doesn’t exist, and (2) the rigid addition operation constrains potential interactions between condition and image tokens.
Instead, OminiControl directly concatenates condition tokens with noisy image tokens $[{X};{C}_{\text{T}};{C}_{\text{I}}]$ for multi-modal attention processing. This formulation enables flexible token interactions through DiT’s multi-modal attention mechanism, allowing direct relationships to emerge between any pair of tokens without imposing rigid spatial constraints.
As shown in Figure 5, this approach effectively handles both spatially aligned and non-aligned tasks, with attention maps revealing clear cross-token relationships and interaction patterns. Empirically, this unified sequence approach consistently achieves lower training loss compared to direct feature addition (Figure 4(a)), demonstrating its superior conditioning capability across diverse generation scenarios.
Figure 5: (a) Attention maps for the Canny-to-image task, showing interactions between noisy image tokens $X$ and image condition tokens $C_{I}$. Strong diagonal patterns indicate effective spatial alignment. (b) Subject-driven generation task, with input condition and output image. Attention maps for $X\to C_{i}$ and $C_{i}\to X$ illustrate accurate subject-focused attention.
Position-aware token interaction
While the unified sequence approach allows flexible token interactions, encoding position information for the newly appended conditional tokens is not straightforward.
In FLUX.1, the RoPE [35] mechanism requires each token to have an assigned position index $(i,j)$, which might directly influence how tokens attend to each other during multi-modal attention.
Specifically, for a 512$\times$512 input image, the VAE encoder produces a 32$\times$32 grid of latent tokens, with each token assigned a position index $(i,j)$ where $i,j\in[0,31]$. Text tokens ${C}_{\text{T}}$ maintain fixed positions at $(0,0)$ following FLUX.1’s convention. Therefore, it is essential to assign proper position indexing the conditioned tokens.
In our method, we assign position indices differently for spatially aligned and non-aligned tasks. For condition image tokens ${C}_{\text{I}}$, we propose a dynamic positioning strategy based on the control task:
$(i,j)_{{C}_{\text{I}}}=\begin{cases}(i,j)_{{X}}&\text{for aligned tasks}\\ (i,j)+\Delta&\text{for non-aligned tasks}\end{cases}$ (4)
where $\Delta$ is a fixed offset vector.
For spatially aligned tasks like edge-guided generation, condition tokens share positions with their corresponding noisy image tokens, facilitating precise spatial correspondence. For non-aligned tasks like subject-driven generation, we shift condition tokens’ positions by $\Delta$, ensuring no spatial overlap with ${X}$. As demonstrated in Figure 4(b), this shifting leads to faster convergence and better performance for non-aligned tasks.
Controllable conditioning strength
While OminiControl’s unified sequence processing and multi-modal attention enable effective token interactions during training, practical applications often require adjustable conditioning strength at test time.
We achieve this by introducing a bias term into the multi-modal attention computation. Specifically, for a given strength factor $\gamma$, we modify the attention operation in Equation 2 to:
$\text{MMA}([{C}_{\text{T}};{X};{C}_{\text{I}}])=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d}}+B(\gamma)\right)V,$ (5)
where $B(\gamma)$ is a bias matrix modulating the attention between concatenated tokens $[{C}_{\text{T}};{X};{C}_{\text{I}}]$. Given ${C}_{\text{T}}\in\mathbb{R}^{M\times d}$ and ${X},{C}_{\text{I}}\in\mathbb{R}^{N\times d}$, the bias matrix has the structure:
$B(\gamma)=\begin{bmatrix}\mathbf{0}_{M\times M}&\mathbf{0}_{M\times N}&\mathbf{0}_{M\times N}\\ \mathbf{0}_{N\times M}&\mathbf{0}_{N\times N}&\log(\gamma)\mathbf{1}_{N\times N}\\ \mathbf{0}_{N\times M}&\log(\gamma)\mathbf{1}_{N\times N}&\mathbf{0}_{N\times N}\end{bmatrix}.$ (6)
This formulation preserves the original attention patterns within each token type while scaling attention weights between ${X}$ and ${C}_{\text{I}}$ by $\log(\gamma)$. At test time, setting $\gamma=0$ removes the condition’s influence, while $\gamma>1$ enhances it. This allows users to flexibly adjust the balance between text and image conditions without retraining the model.
3.3 Subjects200K datasets
Figure 6: Examples from our Subjects200Kdataset. Each pair of images shows the same object in varying positions, angles, and lighting conditions. The dataset includes diverse objects such as clothing, furniture, and animals, totaling over 200,000 images.
A critical challenge in developing universal control frameworks lies in obtaining high-quality training data for subject-driven generation. While OminiControl’s architecture enables flexible control, effective training requires data that maintains subject consistency while incorporating natural variations in pose, lighting, and context.
Existing solutions often rely on identical image pairs [41] or limited-scale datasets [31, 13], which present several limitations. Using identical pairs can lead to overfitting in our framework, causing the model to simply reproduce the input. Meanwhile, existing datasets with natural variations often lack either the scale or diversity needed for robust training.
Condition Model Method Controllability General Quality Text Consistency
[…]
Canny SD1.5 ControlNet 0.34 18.74 0.35 0.45 67.81 0.304
T2I-Adapter 0.22 20.06 0.35 0.39 67.88 0.305
FLUX.1 ControlNet 0.21 98.68 0.25 0.37 56.90 0.192
Ours 0.38 20.63 0.40 0.61 75.91 0.308
Depth SD1.5 ControlNet 923 23.02 0.34 0.47 70.73 0.308
T2I-Adapter 1560 24.72 0.27 0.39 69.99 0.309
FLUX.1 ControlNet 2958 62.20 0.26 0.38 66.84 0.212
Ours 903 27.26 0.39 0.55 75.06 0.308
Deblur FLUX.1 ControlNet 572 30.38 0.74 0.31 54.37 0.293
Ours 132 11.49 0.87 0.39 67.63 0.300
Colorization FLUX.1 ControlNet 351 16.27 0.64 0.43 70.95 0.278
Ours 24 10.23 0.73 0.43 70.74 0.301
Mask SD1.5 ControlNet 7588 13.14 0.40 0.41 67.22 0.300
FLUX.1 Ours 6248 15.66 0.48 0.45 72.61 0.306
Table 1: Quantitative comparison with baseline methods on five spatially aligned tasks. We evaluate methods based on Controllability (F1-Score for Canny, MSE for others), General Quality (FID, SSIM, MAN-IQA, MUSIQ), and Text Consistency (CLIP-Score). For F1-Score, higher is better; for MSE, lower is better. Best results are shown in bold.
To address this data challenge, we leverage a key observation: state-of-the-art DiT models like FLUX.1 [15] can generate identity-consistent image pairs when provided with appropriate prompts. Building on this, we develop an pipeline to create Subjects200K, a large-scale dataset specifically designed for subject-driven generation [^3] Detailed prompt design and synthesization pipeline are provided in the Appendix A.:
  • Prompt Generation. We use GPT-4o to generate over 30,000 diverse subject descriptions. Each description represents the same subject in multiple scenes. The goal is to highlight distinct objects or subjects while allowing natural variations.
  • Paired-image synthesis. We reorganized the collected descriptions into structured prompts. Each prompt describes the same subject in two different scenes. The template of such prompt is shown in Supplementary Material S3.These prompts are then fed into FLUX to generate image pairs. Each pair is designed to maintain subject consistency while varying in pose, lighting, and context.
  • Quality Assesment. Finally, we use GPT-4o to evaluate the generated pairs. Misaligned pairs are removed to ensure identity consistency and high image quality.
The resulting dataset (Figure 6) contains over 200,000 high-quality images spanning diverse categories including clothing, furniture, vehicles, and animals. Each subject appears in multiple contexts, providing rich training signals for learning robust subject-driven control. To facilitate future research, we release both the dataset and our complete generation pipeline [^4] Dataset and code available at: https://github.com/Yuanshi9815/Subjects200K..
4 Experiment
4.1 Setup
Tasks and base model. We evaluate our method on two categories of conditional generation tasks: spatially aligned tasks (including Canny-to-image, depth-to-image, masked-based inpainting, and colorization) and subject-driven generation. We build our method upon FLUX.1 [14], a latent rectified flow transformer model for image generation. By default, we use FLUX.1-dev to generate images for spatially aligned tasks. In subject-driven generation tasks, we switch to FLUX.1-schnell as we observed it tend to produce better visual quality.
Figure 7: Qualitative results comparing different methods. Left: Spatially aligned tasks across Canny, depth, out-painting, deblurring, colorization. Right: Subject-driven generation with beverage can, shoes and robot toy. Our method demonstrates superior controllability and visual quality across all tasks.
Implement details. Our method utilizes LoRA [4]for fine-tuning the base model with a default rank of 4. To preserve the model’s original capabilities and achieve flexibility, the LoRA scale is set to 0 when processing non-condition tokens by default.
Training. Our model is trained with a batch size of 1 and gradient accumulation over 8 steps (effective batch size of 8). We employ the Prodigy optimizer [21] with safeguard warmup and bias correction enabled, setting the weight decay to 0.01. For spatially aligned tasks, we use text-to-image-2M[8] dataset with the last 300,000 images. For subject-driven generation, we utilize our proposed Subjects200K dataset. The experiments are conducted on 2 NVIDIA H100 GPUs (80GB each). For spatially aligned tasks, models are trained for 50,000 iterations, while subject-driven generation models are trained for 15,000 iterations.
Baselines. For spatially aligned tasks, we compare our method with both the original ControlNet [42] and T2I-Adapter [22] on Stable Diffusion 1.5, as well as ControlNetPro [15], the FLUX.1 implementation of ControlNet. For subject-driven generation, we compare with IP-Adapter [41], evaluating its implementations FLUX.1 [39].
Evaluation metrics. We evaluate our model on both spatially aligned tasks and subject-driven generation. For spatially aligned tasks, we assess two aspects: generation quality and controllability. Generation quality is measured using FID [9], SSIM, MAN-IQA [40], and MUSIQ [11] for visual fidelity, along with CLIP Score [27] for semantic consistency. For controllability, we compute F1 Score between extracted and input edge maps in edge-conditioned generation, and MSE between extracted and original condition maps for other tasks (using Depth Anything for depth, color channel separation for colorization, etc.). For subject-driven generation, we propose a five-criteria framework evaluating both preservation of subject characteristics (identity preservation, material quality, color fidelity, natural appearance) and accuracy of requested modifications, with all assessments conducted through GPT-4o’s vision capabilities to ensure systematic evaluation. Detailed evaluation methodology is presented in Appendix B.1.
Evaluation protocol. We conducted evaluations on two datasets. For spatially aligned tasks, we use COCO 2017 validation set (5,000 images) resized to 512$\times$512, using task-specific conditions and associated captions as prompts with a fixed seed of 42. For subject-driven generation, we test on 750 text-condition pairs (30 subjects × 25 prompts) from DreamBooth[31] dataset with 5 different seeds, using one selected image per subject as the condition.
Figure 8: Radar charts visualization comparing our method (blue) with baselines across five evaluation metrics.
4.2 Main result
Spatially aligned tasks As shown in Table 1, we comprehensively evaluate our method against existing approaches on five spatially aligned tasks. Our method achieves the highest F1-Score of 0.38 on depth-to-image generation, significantly outperforming both SD1.5-based methods ControlNet [42] and T2I-Adapter [22], as well as FLUX.1-based ControlNetPro [15]. In terms of general quality metrics, our approach demonstrates consistent superiority across most tasks, showing notably better performance in SSIM [36], MAN-IQA [40], and MUSIQ [11] scores. For challenging tasks like deblurring and colorization, our method achieves substantial improvements: the MSE is reduced by 77% and 93% respectively compared to ControlNetPro, while the FID scores [9] improve from 30.38 to 11.49 for deblurring. The CLIP-Score metrics [27] indicate that our method maintains high text-image consistency across all tasks, suggesting effective preservation of semantic alignment while achieving better control and visual quality. As shown in Figure 7, our method produces sharper details and more faithful color reproduction in colorization tasks, while maintaining better structural fidelity in edge-guided generation and deblurring scenarios.
Methods Base model Parameters Ratio
ControlNet SD1.5 / 860M 361M $\sim$42%
T2I-Adapter 77M $\sim$9.0%
IP-Adapter 449M $\sim$52.2%
ControlNet FLUX.1 / 12B 3.3B $\sim$27.5%
IP-Adapter 918M $\sim$7.6%
Ours FLUX.1 / 12B 14.5M / 48.7M w/ Encoder $\sim$0.1/ $\sim$0.4w/ Encoder
Table 2: Additional parameters introduced by different image conditioning methods. For IP-Adapter, the parameter count includes the CLIP Image encoder. For our method, we also report results when using the original VAE encoder from FLUX.1.
Subject driven generation Figure 8 presents a comprehensive comparison against existing baselines. Our method demonstrates superior performance, particularly in identity preservation and modification accuracy. Averaging over random seeds, we achieve 75.8% modification accuracy compared to IP-Adapter (FLUX)’s 57.7%, while maintaining 50.6% identity preservation against IP-Adapter (SD 1.5)’s 29.4%. The advantage amplifies in best-seed scenarios, achieving 90.7% modification accuracy and 82.3% identity preservation - surpassing the strongest baselines by 15.8 and 18.0 percentage points, demonstrating effective subject-fidelity editing. These quantitative results are further corroborated by user studies presented in Appendix B.1.
Comparative paramter efficiency. As shown in Table 2, our approach achieves remarkable parameter efficiency compared to existing methods. For the 12B parameter FLUX.1 model, our method requires only 14.5M trainable parameters (approximately 0.1%), which is significantly lower than ControlNet (27.5%) and IP-Adapter (7.6%). Even when utilizing the original VAE encoder from FLUX.1, our method still maintains high efficiency with just 0.4% additional parameters, demonstrating the effectiveness of our parameter-efficient design.
4.3 Empirical studies
Effect of training data.
Figure 9: Comparison of models trained with different data. The model trained by data augmentation tends to copy inputs directly, while model trained by our Subjects200Kgenerates novel views while preserving identity.
For subject-driven generation, our model takes a reference image of a subject (e.g., a plush toy or an object) and a text description as input, aiming to generate novel images of the same subject following the text guidance while preserving its key characteristics.
To validate the effectiveness of our Subjects200K dataset described in Section 3.3, we compare two training strategies for this task. The first approach relies on traditional data augmentation, where we apply random cropping, rotation, scaling, and adjustments to contrast, saturation, and color to the original images. The second approach utilizes our Subjects200K dataset. As shown in Figure 9, the model trained with data augmentation only learns to replicate the input conditions with minimal changes. In the first row, it simply places the taco plush toy in a bright room setting while maintaining its exact appearance and pose. Similarly, in the second row, the yellow alarm clock is reproduced with nearly identical details despite the window-side placement instruction. In contrast, our Subjects200K-trained model demonstrates the ability to generate diverse yet consistent views of the subjects while faithfully following the text prompts.
Condition strength analysis. We evaluate our condition strength control through qualitative experiments. Figure 10 shows generated results with varying strength factor $\gamma\in{0.25,0.5,0.75,1.0}$. Results show that $\gamma$ effectively controls the generation process for both spatially aligned tasks like depth-to-image generation and spatially non-aligned tasks like subject-driven generation, enabling flexible control over the condition’s influence.
Figure 10: Comparison of models trained with different data. The model trained by data augmentation tends to copy inputs directly, while model trained by our Subjects200Kgenerates novel views while preserving identity.
Impact of LoRA rank. We conducted extensive experiments with different LoRA ranks (1, 2, 4, 8, and 16) for the Canny-to-image task. As shown in Table 3, our experiments show that increasing the LoRA rank generally improves model performance, with rank 16 achieving the best results across multiple aspects: image quality (measured by FID and SSIM), condition control capability (measured by F1 Score), while maintaining competitive text-image consistency (measured by CLIP-Score). Notably, even with smaller ranks (e.g., rank 1), the model demonstrates competitive performance, especially in text-image alignment where it achieves the highest CLIP-Score of 0.765, showing the efficiency of our approach even with limited parameters.
Study Setting FID $\downarrow$ SSIM $\uparrow$ F1 Score $\uparrow$ CLIP Score $\uparrow$
LoRA Rank 1 21.09 0.412 0.385 0.765
2 21.28 0.411 0.377 0.751
4 20.63 0.407 0.380 0.761
8 21.40 0.404 0.3881 0.761
16 19.71 0.425 0.407 0.764
Condition Blocks Early 25.66 0.369 0.23 0.72
Full 20.63 0.407 0.38 0.76
Table 3: Ablation studies on (1) LoRA rank for the Canny-to-image task and (2) condition signal integration approaches. Results show that LoRA rank of 16 and full-depth integration achieve the best performance. Rows with blue background indicate our default settings (LoRA rank=4, Full condition integration). Best results are in bold.
Conditioning depth. FLUX.1’s transformer architecture features two distinct types of blocks: early blocks that employ separate normalization modules for different modalities tokens (text and image) and later blocks that share unified normalization across all tokens. As shown in Table 3, experiments reveal that restricting condition signal integration to only these early blocks results in insufficient controllability over the generation process. This suggests that allowing the condition signals to influence the entire transformer stack is crucial for achieving the desired levels of control over the output. Notably, this finding indicates that the preview approaches[15, 39, 41, 42, 22] of inserting condition signals primarily in early blocks, which were effective in UNet-based architectures, may not fully translate to DiT-based models like FLUX.1.
5 Conclusion
OminiControl offers parameter-efficient image-conditioned control for Diffusion Transformers across diverse tasks using a unified token approach, without extra modules. Our method outperforms traditional approaches, and the new Subjects200K dataset—featuring over 200,000 high-quality, subject-consistent images—supports advancements in subject-consistent generation. Results confirm OminiControl’s scalability and effectiveness in diffusion models.
6 Acknowledgment
We would like to acknowledge that the computational work involved in this research work is partially supported by NUS IT’s Research Computing group using grant numbers NUSREC-HPC-00001.
Supplementary Material
Appendix A Details of Subjects200K datasets
We present a comprehensive synthetic dataset constructed to address the limitations in scale and image quality found in previous datasets [31, 13, 16, 17]. Our approach leverages FLUX.1-dev [14] to generate high-quality, consistent images of the same subject under various conditions.
Subjects200K dataset currently consists of two splits, both generated using similar pipelines. Split-1 contains paired images of objects in different scenes, while Split-2 pairs each object’s scene image with its corresponding studio photograph. Due to their methodological similarities, we primarily focus on describing the synthesis process and details of Split-2, although both splits are publicly available. Our complete Subjects200K dataset can be fully accessed via this link.
A.1 Generation pipeline
Our dataset generation process consists of three main stages: description generation, image synthesis, and quality assessment.
Description Generation We employed ChatGPT-4o to create a hierarchical structure of descriptions: We first generated 42 diverse object categories, including furniture, vehicles, electronics, clothing, and others. For each category, we created multiple object instances, totaling 4,696 unique objects. Each object entry consists of: (1) A brief description, (2) Eight diverse scene descriptions, (3) One studio photo description. Figure S2 shows a representative example of our structured description format.
Image Synthesis We designed a prompt template to leverage FLUX’s capability of generating paired images containing the same subject. Our template synthesizes a comprehensive prompt by combining a brief object description with two distinct scene descriptions, ensuring subject consistency while introducing environmental variations.
The detailed prompt structure is illustrated in Figure S3. For each prompt, we set the image dimensions to 1056×528 pixels and generated five images using different random seeds to ensure diversity in our dataset. During the training process, we first split the paired images horizontally, then performed central cropping to obtain 512×512 pixel image pairs. This padding strategy was implemented to address cases where the generated images were not precisely bisected, preventing potential artifacts from appearing in the wrong half of the split images.
Figure S1: Examples of successful and failed generation results from Subjects200K dataset. Green checks indicate successful cases where subject identity and characteristics are well preserved, while red crosses show failure cases.
Quality assessment We leveraged ChatGPT-4o’s vision capabilities to rigorously evaluate the quality of images generated by FLUX.1-dev. The assessment focused on multiple critical aspects:
  • Image composition: Verifying that each image properly contains two side-by-side views.
  • Subject consistency: Ensuring the subject maintains identity across both views.
  • Image quality: Confirming high resolution and visual fidelity.
To maintain stringent quality standards, each image underwent five independent evaluations by ChatGPT-4o. Only images that passed all five evaluations were included in our training dataset. Figure S1 presents representative examples from our quality-controlled dataset.
"brief_description":
"A finely-crafted wooden seating piece.",
"scene_descriptions": [
"Set on a sandy shore at dusk, it faces the ocean with a gentle breeze rustling nearby palms, bathed in soft, warm twilight.",
"Positioned in a bustling urban cafe, it stands out against exposed brick walls, capturing the midday sun through a wide bay window."
// Additional six scene descriptions omitted
"studio_photo_description":
"In a professional studio against a plain white backdrop, it is captured in three-quarter view under uniform high-key lighting, showcasing the delicate grain and smooth of its finely-crafted surfaces."
Figure S2: An example of our structured description format for dataset generation.
prompt_1 = f"Two side-by-side images of the same object: {brief_description}"
prompt_2 = f"Left: {scene_description1}"
prompt_3 = f"Right: {scene_description2}"
prompt_image = f"{prompt_1}; {prompt_2}; {prompt_3}"
Figure S3: Our prompt template for paired image generation. The template combines a brief object description with two distinct scene descriptions to maintain subject consistency while varying environmental conditions.
A.2 Dataset Statistics
In Split-2, we first generated 42 distinct object categories, from which we created and curated a set of 4,696 detailed object instances. Then we combine these descriptions to generate 211,320 subject-consistent image pairs. Through rigorous quality control using GPT-4o, we selected 111,767 high-quality image pairs for our final dataset. This extensive filtering process ensured the highest standards of image quality and subject consistency, resulting in a collection of 223,534 high-quality training images.
Appendix B Additional experimental results
B.1 Evaluation for subject-driven generation
Framework and criteria. To systematically evaluate subject-driven generation quality, we establish a framework with five criteria assessing both preservation of subject characteristics and accuracy of requested modifications:
  • Identity Preservation: Evaluates preservation of essential identifying features (e.g., logos, brand marks, distinctive patterns)
  • Material Quality: Assesses if material properties and surface characteristics are accurately represented
  • Color Fidelity: Evaluates if colors remain consistent in regions not specified for modification
  • Natural Appearance: Assesses if the generated image appears realistic and coherent
  • Modification Accuracy: Verifies if the changes specified in the text prompt are properly executed
Method Identity Material Color Natural Modification Average
preservation quality fidelity appearance accuracy score
Average over 5 random seeds
IP-Adapter (SD 1.5) 29.4 86.1 45.3 97.9 17.0 55.1
SSR-Encoder 46.0 92.0 54.2 96.3 28.5 63.4
IP-Adapter (FLUX) 11.8 65.8 30.8 98.1 57.7 52.8
Ours 50.6 84.3 55.0 98.5 75.8 72.8
Best score over 5 random seeds
IP-Adapter (SD 1.5) 56.3 98.9 70.1 99.7 37.2 72.5
SSR-Encoder 64.3 99.2 74.4 99.1 53.6 78.1
IP-Adapter (FLUX) 27.5 86.1 53.6 99.9 74.9 68.4
Ours 82.3 98.0 88.4 100.0 90.7 91.9
Table S1: Quantitative evaluation results (in percentage) across different evaluation criteria. Higher values indicate better performance.
Figure S4: User study results comparing different methods across three metrics: identity consistency, text-image alignment, and visual coherence.
User studies. To further validate our approach, we conducted user studies collecting 375 valid responses. Participants evaluated the generated images across three key dimensions: identity consistency, text-image alignment, and visual coherence between subjects and backgrounds. The results shown in Figure S4 corroborate our quantitative findings, with our method achieving superior performance across all evaluation criteria.
B.2 Additional generation results
We showcase more generation results from our method. Figure S5 presents additional results on the DreamBooth dataset, while Figure S6 demonstrates our method’s effectiveness on other subject-driven generation tasks.
Figure S5: More results on Dreambooth dataset.
Figure S6: More results on other subject-driven generation tasks.